OpenStudy (anonymous):

Alex was finding the average change of his stocks over the following months: -10.29 in June, -8.16 in July, 34.50 in August, 3.12 in September. Estimate the average change in his stocks by rounding first to the nearest whole number and then dividing by the number of months.

OpenStudy (anonymous):

Pretty much asking you to estimate the difference quotient: \[\frac{f(x+h)-f(x)}{h}\] June to July: \[\frac{f(7)-f(6)}{7-6} \approx \frac{10-(-8)}{1} \approx 18\] Follow similar steps for June to August, July to August

OpenStudy (anonymous):

and september